OpenStudy (shadowcain):

x ---- - 8 < -2 -2 ---

OpenStudy (calculusxy):

\[\frac{ x }{ -2 } - 8 < -2\] is that your question?

OpenStudy (calculusxy):

if that's so \[\frac{ x }{ -2 } - 8 < -2\] \[\frac{ x }{ -2 } -8 + 8 < -2 + 8\] \[\frac{ x }{ -2 } < 6\] \[x < 6 \times -2\] \[x < -12\]

OpenStudy (shadowcain):

So, it would be B

OpenStudy (calculusxy):

no what B represents is \[x \le -12 \] since the circle is shaded instead it would be A because the circle is unshaded

OpenStudy (calculusxy):

@ShadowCain

OpenStudy (calculusxy):

\[5 < \frac{ x }{ -3 } + 1 \] \[4 < \frac{ x }{ -3 }\] \[-12 < x \]